# Comparing Quantum Annealing and Spiking Neuromorphic Computing for Sampling Binary Sparse Coding QUBO Problems

### v1

@@ Abstract
We consider the problem of computing a sparse binary representation of an image. To be precise, given an image and an overcomplete, non-orthonormal basis, we aim to find a sparse binary vector indicating the minimal set of basis vectors that when added together best reconstruct the given input. We formulate this problem with an $L_{2}$ loss on the reconstruction error, and an $L_{0}$ (or, equivalently, an $L_{1}$) loss on the binary vector enforcing sparsity. This yields a quadratic binary optimization problem (QUBO), whose optimal solution(s) in general is NP-hard to find. The contribution of this work is twofold. First, the method of unsupervised and unnormalized dictionary feature learning for a desired sparsity level to best match the data is presented. Second, we solve the sparse representation QUBO by implementing it both on a D-Wave quantum annealer with Pegasus chip connectivity via minor embedding, as well as on the Intel Loihi 2 spiking neuromorphic processor. On the D-Wave quantum annealer, we sample from the sparse representation QUBO using parallel quantum annealing combined with quantum evolution Monte Carlo, also known as iterated reverse annealing. On the Loihi 2 neuromorphic chip, we use a stochastic winner take all network of neurons. The solutions are benchmarked against simulated annealing, a classical heuristic, and the optimal solutions are computed using CPLEX. Iterated reverse quantum annealing performs similarly to simulated annealing, although simulated annealing is always able to sample the optimal solution whereas quantum annealing was not always able to. The Loihi 2 solutions that are sampled are on average more sparse than the solutions from any of the other methods. We demonstrate that both quantum annealing and neuromorphic computing are suitable for binary sparse coding QUBOs, and that Loihi 2 outperforms a D-Wave quantum annealer standard linear-schedule anneal, while for these types of QUBOs, iterated reverse quantum annealing performs much better than both unmodified linear-schedule quantum annealing and iterated warm starting on Loihi 2.
@@ 1 Introduction
Mathematically, we are given an overcomplete and non-orthonormal basis of $n\in\mathbb{N}$ matrices $\bm{\phi}=\{\phi_{1},\ldots,\phi_{n}\}$, and we aim to infer a sparse representation of a given image $\bm{I}$. Here, an overcomplete set is defined as one that contains more functions than needed for a basis. All basis matrices as well as the image $\bm{I}$ are assumed to be of equal dimensions. The task is to find the minimal set of non-zero activation coefficients $\bm{a}$ that accurately reconstruct $\bm{I}$, where $\bm{a}\in\mathbb{B}^{n}$ is a binary vector of length $n$ for $\mathbb{B}=\{0,1\}$. We can express the computation of a sparse binary representation of the image $\bm{I}$ using the basis $\bm{\phi}$ as the minimization of the energy function
$E\left(\bm{I},\bm{a}\right)=\min_{\bm{a}}\left[\frac{1}{2}\|\bm{I}-\bm{\phi}\bm{a}\|_{2}^{2}+\lambda\|\bm{a}\|_{0}\right],$ (1)
@@
where $\|\cdot\|_{2}$ is the Euclidean norm and $\|\cdot\|_{0}$ denotes the number of nonzero elements. The operation $\bm{\phi}\bm{a}$ denotes the addition of all basis matrices in $\bm{\phi}$ that correspond to an entry of $1$ in $\bm{a}$. The parameter $\lambda>0$ is a Lasso-type penalization parameter controlling the sparseness of the solution [Tibshirani1996]. A large value of $\lambda$ results in a more sparse solution to eq. (1), while smaller values yield denser solutions. Therefore, the parameter $\lambda$ allows one to effectively balance the reconstruction error (the $L_{2}$ norm) and the number of non-zero activation coefficients (the $L_{0}$ norm). Since eq. (1) belongs to the class of 0-1 integer programming problems, finding a sparse representation falls into an NP-hard complexity class. The energy function of eq. (1) is non-convex and typically contains multiple local minima.
@@
We investigate two types of computing models to solve the sparse representation problem given in eq. (1). The first is a D-Wave quantum annealer with chip ID Advantage_system4.1. Moreover, we employ the spiking neuromorphic computing processor Loihi 2 by Intel. Quantum annealing is a type of analog quantum computation that aims to use quantum fluctuations in order to search the objective function landscape of a combinatorial optimization problem for the global minimum [Kadowaki_1998, Farhi2000]. Quantum annealer devices are manufactured by D-Wave, which use superconducting flux qubits to implement quantum annealing with a transverse field Hamiltonian and have been studied for sampling a variety of problem types [Johnson2011, Lanting_2014]. Although current quantum annealers are subject to various sources of errors, limited coherence times, and limited hardware connectivity, there is evidence that quantum annealing can provide an advantage over classical algorithms for certain types of problems [King2021scaling, Tasseff2022, king2024computational, bauza2024scaling]. Neuromorphic computing is a proposed computing model inspired by the human brain, which is able to complete certain learning tasks better than classical von Neumann computers [Davies2021, Schuman2022].
@@
This article is a significant extension of [henke2023sampling], in which the unsupervised dictionary learning approach to generate sparse coding QUBO models was introduced for the purposes of sampling those problems on Loihi 1. The article is structured as follows. After a literature review in Section 1.1, we introduce the quantum annealing and neuromorphic computing technology we investigate in Section 2. Experimental results are presented in Section 3. The article concludes with a discussion in Section 4.
@@ 1.1 Literature review
The goal of any sparse coding method is to accurately recover an unknown sparse vector from a few noisy linear measurements. Unfortunately, this estimation problem is NP-hard in general (when formulated as 0-1 integer optimization), and it is therefore usually approached with an approximation method, such as some Lasso or orthogonal matching pursuit (OMP) [4385788], thus trading off accuracy for less computational complexity.
In [Rozell2008], the authors investigate the suitability of sparse approximations, or also called sparse coding, to minimize an objective consisting of an error term with $L_{2}$ loss and a coefficient cost function, where sparse representations are enforced in neural systems with the help of local symmetric competitions among neurons.
A comparison of quantum computing (using the D-Wave quantum annealer) and neuromorphic computing (using the Intel Loihi spiking processor) has been investigated in [Henke2020machine, Henke2020alien], who compare both on an identically-posed machine learning problem with respect to various metrics (power consumption, reconstruction and classification accuracy, etc). This line of research is continued in [Henke2022], who compare sparse coding solutions generated classically by orthogonal matching pursuit (OMP) with spike-based solutions obtained with the Loihi 1 neuromorphic processor. Their article investigates post-hoc normalization techniques to speed up computations on Loihi 1, optimal parameter selection, reconstruction errors, and unsupervised dictionary learning for Loihi 1 approaches and their classical counterparts.
A quantum-inspired algorithm for sparse coding is presented in [Romano2022] by formulating a general sparse coding problem as a quadratic unconstrained binary optimization (QUBO) problem suitable for approximation on current hardware such as the D-Wave quantum annealers. Similar research on regularized linear regression for sparse signal reconstruction is presented in [Ide2022], who also derive a QUBO representation and compare solutions obtained with D-Wave Advantage Systems 4.1 to OMP and conventional Lasso methods.
Two related problems, binary compressive sensing and binary compressive sensing with matrix uncertainty, are investigated in [Ayanzadeh2019], though no experimental results on D-Wave devices are presented.
@@ 2 Methods
In this section we present the hardware being used in this research. After some mathematical considerations on the problem under investigation (Section 2.1), we give details on Intel’s Loihi 2 neuromorphic chip (Section 2.2), followed by some implementation details for the D-Wave quantum annealer (Section 2.3). Section 2.4 describes the classical simulated annealing implementation which we will compare against. Section 2.5 discusses the exact solution obtained with IBM’s CPLEX solver [CPLEX], which is exact and deterministic.
@@ 2.1 Binary Sparse Coding QUBOs
We start by reformulating eq. (1) in QUBO form. To this end, denoting $D=\bm{\phi}$ and $\bm{x}=\bm{I}$, we observe that for $\bm{a}\in\mathbb{B}^{n}$,
@@
$=\frac{1}{2}\bm{x}^{\top}\bm{x}-\bm{x}^{\top}D\bm{a}+\frac{1}{2}\bm{a}^{\top}D^{\top}D\bm{a}+\lambda\sum_{i=1}^{n}a_{i},$
where $\bm{a}=(a_{1},\ldots,a_{n})$. As expected, multiplying out eq. (1) yields a quadratic form in $\bm{a}$, meaning that we can recast our objective function as a QUBO problem. For this we define the following two transformations:
$h_{i}$ $=-D_{i}^{\top}\bm{x}+\lambda+\frac{1}{2}D_{i}^{\top}D_{i}$ (2)
$Q_{ij}$ $=\frac{1}{2}(D^{\top}D)_{i}-\mathbb{I}(i=j),$ (3)
where $\mathbb{I}$ denotes the indicator function which returns $1$ if and only if its argument is true, and 00 otherwise. Using eq. (2) and eq. (3), we can rewrite eq. (1) as the following QUBO,
$H(\bm{h},Q,\bm{a})=\sum_{i=1}^{n}{h_{i}a_{i}}+\sum_{i<j}Q_{ij}a_{i}a_{j}$ (4)
@@
Figure 1: Colormaps of the full QUBO symmetric matrix for a (randomly chosen and representative) sparse coding QUBO with 64 variables (left) and 128 variables (right). Each heatmap corresponds to the coefficient for each of the linear terms (on the diagonal) and for each of the quadratic terms (on the off-diagonals).
@@
The binary sparse coding QUBOs are constructed using the un-normalized dictionary learning introduced in [henke2023sampling]. In particular, the dictionary learning is performed using the classical heuristic simulated annealing using the settings that will be described in Section 2.4. As done previously [henke2023sampling], a single image from the standard fashion MNIST (fMNIST) data set [xiao2017fashionmnist] is used, where the 28x28 image is divided into 16 7x7 patches, and then a dictionary size of $64$ variables is used for each of the $16$ patches. We also evaluate a version where we increase the dictionary size to $128$ variables, and use a patch size of 2x2 (resulting in $4$ separate QUBOs).
2.2 Implementation on Loihi

### v2

@@ Abstract
We consider the problem of computing a sparse binary representation of an image. To be precise, given an image and an overcomplete, non-orthonormal basis, we aim to find a sparse binary vector indicating the minimal set of basis vectors that when added together best reconstruct the given input. We formulate this problem with an $L_{2}$ loss on the reconstruction error, and an $L_{0}$ (or, equivalently, an $L_{1}$) loss on the binary vector enforcing sparsity. This yields a quadratic unconstrained binary optimization problem (QUBO), whose optimal solution(s) in general is NP-hard to find. The contribution of this work is twofold. First, we solve the sparse representation QUBOs by solving them both on a D-Wave quantum annealer with Pegasus chip connectivity via minor embedding, as well as on the Intel Loihi 2 spiking neuromorphic processor using a stochastic Non-equilibrium Boltzmann Machine (NEBM). Second, we deploy Quantum Evolution Monte Carlo with Reverse Annealing and iterated warm starting on Loihi 2 to evolve the solution quality from the respective machines. The solutions are benchmarked against simulated annealing, a classical heuristic, and the optimal solutions are computed using CPLEX. Iterated reverse quantum annealing performs similarly to simulated annealing, although simulated annealing is always able to sample the optimal solution whereas quantum annealing was not always able to. The Loihi 2 solutions that are sampled are on average more sparse than the solutions from any of the other methods. We demonstrate that both quantum annealing and neuromorphic computing are suitable for binary sparse coding QUBOs, and that Loihi 2 outperforms a D-Wave quantum annealer standard linear-schedule anneal, while iterated reverse quantum annealing performs much better than both unmodified linear-schedule quantum annealing and iterated warm starting on Loihi 2.
@@ 1 Introduction
Mathematically, we are given an overcomplete and non-orthonormal basis of $p\in\mathbb{N}$ vectors $D=\{D_{1},\ldots,D_{p}\}$, and we aim to infer a sparse representation of a given image $\bm{x}$. Here, an overcomplete set is defined as one that contains more functions than needed for a basis. All basis vectors and the image $\bm{x}$ are assumed to be of equal dimension $m\in\mathbb{N}$. The task is to find the minimal set of non-zero activation coefficients $\bm{a}$ that accurately reconstruct $\bm{x}$, where $\bm{a}\in\mathbb{B}^{p}$ is a binary vector of length $p$ for $\mathbb{B}=\{0,1\}$. We can express the computation of a sparse binary representation of the image $\bm{x}$ using the basis $D$ as the minimization of the energy function
$E\left(\bm{a}\right)=\min_{\bm{a}}\left[\frac{1}{2}\|\bm{x}-D\bm{a}\|_{2}^{2}+\lambda\|\bm{a}\|_{0}\right],$ (1)
@@
where $\|\cdot\|_{2}$ is the Euclidean norm and $\|\cdot\|_{0}$ denotes the number of nonzero elements. The operation $D\bm{a}$ denotes the addition of all basis matrices in $D$ that correspond to an entry of $1$ in $\bm{a}$. The parameter $\lambda>0$ is a Lasso-type penalization parameter controlling the sparseness of the solution [Tibshirani1996]. A large value of $\lambda$ results in a more sparse solution to eq. (1), while smaller values yield denser solutions. Therefore, the parameter $\lambda$ allows one to effectively balance the reconstruction error (the $L_{2}$ norm) and the number of non-zero activation coefficients (the $L_{0}$ norm). Since eq. (1) belongs to the class of 0-1 integer programming problems, finding a sparse representation falls into an NP-hard complexity class. The energy function of eq. (1) is non-convex and typically contains multiple local minima.
@@
We investigate two types of computing models to solve the sparse representation problem given in eq. (1). The first is a D-Wave quantum annealer with chip ID Advantage_system4.1. Moreover, we employ the spiking neuromorphic computing processor Loihi 2 by Intel. Quantum annealing is a type of analog quantum computation that aims to use quantum fluctuations in order to search the objective function landscape of a combinatorial optimization problem for the global minimum [Kadowaki_1998, Farhi2000]. Quantum annealer devices are manufactured by D-Wave, which use superconducting flux qubits to implement quantum annealing with a Transverse field Hamiltonian and have been studied for sampling a variety of problem types [Johnson2011, Lanting_2014]. The Transverse field driving Hamiltonian facilitates state transitions. Although current quantum annealers are subject to various sources of errors, limited coherence times, and limited hardware connectivity, there is evidence that quantum annealing can provide an advantage over classical algorithms for certain types of problems [King2021scaling, Tasseff2022, king2024computational, bauza2024scaling]. Neuromorphic computing is a proposed computing model inspired by the human brain, which is able to complete certain learning tasks better than classical von Neumann computers [Davies2021, Schuman2022]. To the best of our knowledge, this is the first demonstration of an exact $L_{0}$ constraint problem being solved on two the two computing paradigms, respectively.
@@
This article is a significant extension of [henke2023sampling], in which the unsupervised dictionary learning approach to generate sparse coding QUBO models was introduced for the purposes of sampling those problems on Loihi 1. Among others, this article investigates post-hoc normalization techniques to speed up computations on Loihi 1, optimal parameter selection, reconstruction errors, and unsupervised dictionary learning for Loihi 1 approaches and their classical counterparts. The article is structured as follows. After a literature review in Section 1.1, we introduce the quantum annealing and neuromorphic computing technology we investigate in Section 2. Experimental results are presented in Section 3. The article concludes with a discussion in Section 4. Data generated by this study is publicly available [pelofske_2025_14735999].
@@ 1.1 Literature Review
The goal of any sparse coding method is to accurately recover an unknown sparse vector from a few noisy linear measurements. Unfortunately, this estimation problem is NP-hard in general (when formulated as 0-1 integer optimization), and it is therefore usually approached with an approximation method, such as some (C)LASSO (Constrained Least Absolute Shrinkage and Selection Operator) or orthogonal matching pursuit (OMP) [4385788] where the penalty term in eq. (1) is an $L_{1}$ ($\left\|\cdot\right\|_{1}$) constraint. Because the coefficients are allowed to be any positive real number, the problem is turned convex, and a global minimum is achieved at the cost of trading off accuracy for less computational complexity.
A comparison of quantum computing (using a D-Wave quantum annealer) and neuromorphic computing (using the Intel Loihi spiking processor) has been investigated in [Henke2020machine, Henke2020alien]. The binary results of the D-Wave quantum annealer and the rate coded solutions motivated by [Rozell2008] to the $L_{1}$ problem from Loihi 1 were examined with respect to various metrics (power consumption, reconstruction and classification accuracy, etc), which generally finds both the quantum annealing and neuromorphic processors to be reasonably effective at solving the respective underlying problem – however, exact state-of-the-art classical algorithms outperform both methods. This line of research is continued in [Henke2022], who compare sparse coding solutions generated classically by a greedy orthogonal matching pursuit (OMP) algorithm with spike-based solutions obtained with the Loihi 1 neuromorphic processor.
A quantum-inspired algorithm for sparse coding is presented in [Nguyen_2018, Romano2022] by formulating a general sparse coding problem as a quadratic unconstrained binary optimization (QUBO) problem suitable for approximation on current hardware such as D-Wave quantum annealers. The QUBO formulations of [Nguyen_2018, Romano2022] were shown to be competitive with the standard (C)LASSO and OMP methods. Similar research on regularized linear regression for sparse signal reconstruction is presented in [Ide2022], who also derive a QUBO representation and compare solutions obtained with D-Wave Advantage Systems 4.1 to OMP and conventional (C)LASSO methods, the results for which showed good recovery success rates using the quantum annealer that is comparable to conventional methods.
Two related problems, binary compressive sensing and binary compressive sensing with matrix uncertainty, are investigated in [Ayanzadeh2019]; this study presents an Ising model formulation that has a ground state that is a sparse solution for the binary compressive sensing problem, though no experimental results on D-Wave devices are presented.
@@ 2 Methods
In this section we present the hardware being used in this research. After some mathematical considerations on the problem under investigation (Section 2.1), we give details on Intel’s Loihi 2 neuromorphic chip (Section 2.3.6), followed by some implementation details for the D-Wave quantum annealer (Section 2.4). Section 2.5 describes the classical simulated annealing implementation which we will compare against. Section 2.6 discusses the exact solution obtained with IBM’s CPLEX solver [CPLEX], which is exact and deterministic.
@@ 2.1 Binary Sparse Coding QUBOs
We start by reformulating eq. (1) in QUBO form. First we expand out the objective function,
@@
$=\frac{1}{2}\left[(\bm{x}-D\bm{a})^{\top}(\bm{x}-D\bm{a})\right]+\lambda\sum_{i=1}^{p}a_{i}$
$=\frac{1}{2}(\bm{x}^{\top}\bm{x}-2\bm{x}^{\top}D\bm{a}+\bm{a}^{\top}D^{\top}D\bm{a})+\lambda\sum_{i=1}^{p}a_{i}$
where $\bm{a}=(a_{1},\ldots,a_{p})$. Next, we take the derivative with respect to one component of $\bm{a}$,
$\begin{split}\partial_{a_{i}}E(\bm{a})={}&\partial_{a_{i}}\left[\frac{1}{2}(\sum_{j=1}^{m}x_{j}^{2}-2\sum_{j=1}^{m}(x_{j}\sum_{i=1}^{p}D_{ji}a_{i})+\sum_{j=1}^{m}(\sum_{i=1}^{p}D_{ji}a_{i})^{2})+\lambda\sum_{i=1}^{p}a_{i}\right]\end{split}$
$\begin{split}={}&-\sum_{j=1}^{m}x_{j}D_{ji}+\sum_{j=1}^{m}(\sum_{i=1}^{p}D_{ji}a_{i})D_{ji}+\lambda\end{split}$
$\begin{split}={}&-\bm{x}^{\top}D_{i}+D_{i}^{\top}D\bm{a}+\lambda.\end{split}$ (2)
As expected, multiplying out eq. (1) yields a quadratic form in $\bm{a}$, and taking the derivative yields an expression in terms of $\bm{a}$. Hence we can recast our objective function as a QUBO problem. For this we define the following two transformations:
$h_{i}$ $=-\bm{x}^{\top}D_{i}+D_{i}^{\top}D_{i}+\lambda,$ (3)
$Q_{i}$ $=D_{i}^{\top}D.$ (4)
Using eq. (3) and eq. (4), where $D_{i}^{\top}D_{i}$ represents the self interaction term and hence can be absorbed into the linear term, we can rewrite eq. (1) as the following QUBO,
$H(\bm{h},Q,\bm{a})=\sum_{i=1}^{p}{h_{i}a_{i}}+\sum_{i<j}Q_{ij}a_{i}a_{j},$ (5)
@@
Figure 1: Binary sparse coding QUBO model coefficient matrices. Colormaps of the full QUBO symmetric matrix for a (randomly chosen and representative) sparse coding QUBO with 64 variables (left) and 128 variables (right). Each heatmap corresponds to the coefficient for each of the linear terms (on the diagonal) and for each of the quadratic terms (on the off-diagonals). The axis labels correspond to the matrix row and column indices.
@@
The binary sparse coding QUBOs are constructed using the un-normalized dictionary learning introduced in [henke2023sampling]. In particular, the dictionary learning is performed using the classical heuristic simulated annealing using the settings that will be described in Section 2.5. As done previously [henke2023sampling], a single image from the standard fashion MNIST (fMNIST) data set [xiao2017fashionmnist] is used, where the 28x28 image is divided into 16 7x7 patches, and then a dictionary size of $64$ variables is used for each of the $16$ patches. We also evaluate a version where we increase the dictionary size to $128$ variables.
2.2 Unsupervised Dictionary Learning
Sparse coding optimization can be viewed as a two-step process. First, a dictionary of basis vectors is learned in an unsupervised manner using a local Hebbian rule. When solving the convex (C)LASSO problem, it is typically necessary to re-normalize the columns of the dictionary $D$ after each learning epoch. This normalization is essential for convergence in (C)LASSO because the entries of the sparse vector $a$ can take on any real value.
Here, we use a learning procedure presented in [henke2023sampling] that enables the algorithm to determine the optimal feature norms based on a specified target average sparsity level, $s\in(0,1)$. The dictionary is initialized with features drawn from a normal distribution, ensuring their norms are randomly below 1, along with a small sparsity penalty parameter $\lambda$. After solving the binary sparse coding problem for each sample in the training set, the dictionary is updated. If the average sparsity across the training epoch exceeds the target $s$, the penalty parameter $\lambda$ is increased for the next epoch until the average reconstruction error and sparsity converge.
2.3 Loihi 2 Neuromorphic Chip Implementation
The Intel Loihi 2 neuromorphic processor is designed to replicate core elements of biological neural structure and processing, thereby targeting specific classes of computational tasks. Compared to its predecessor, Loihi 1, Loihi 2 offers several enhancements, including improved manufacturing processes, more connectivity, faster processing, and lower energy consumption. Each Loihi 2 chip contains roughly one million neurons and approximately 120 million synapses.
Loihi hardware employs binary spike events, wherein neurons communicate through digital pulses that can be either “on” or “off” (1 or 0). This binary scheme mirrors the all-or-nothing firing patterns observed in biological neurons, though with lower granularity.
Whereas the Loihi 1 architecture has been utilized with the Stochastic Constraint Integrate and Fire (SCIF) neuron model [Aimone_2022, henke2023sampling] to solve QUBO problems, the expanded capabilities of Loihi 2 support more sophisticated models. One such model is the Non-equilibrium Boltzmann Machine (NEBM) neuron model, available in Intel’s lava software package and specifically designed for optimization tasks such as Quadratic Unconstrained Binary Optimization (QUBO). Drawing on principles of statistical mechanics and thermodynamics—particularly the concept of a non-equilibrium steady state [PhysRevLett.112.180602]—the NEBM model departs from traditional approaches by operating under continuously driven conditions rather than settling into thermal equilibrium.
In contrast to equilibrium-based Boltzmann machines, which treat neurons as binary units obeying a Boltzmann distribution, NEBM introduces dynamics influenced by external driving forces. This design encourages more extensive exploration of the solution space and offers a mechanism to escape local minima. Additionally, NEBM does not conserve energy in the traditional sense; it instead dissipates energy due to the open, driven nature of the system. This dissipation is essential for creating an effective annealing process, wherein the system is gradually “cooled” to stabilize into low-energy configurations that correspond to high-quality solutions. Unlike strictly equilibrium-based models, NEBM inherently incorporates time-dependent behavior, allowing for dynamic update rules that can be tuned for faster convergence or more thorough exploration.
When applying NEBM to a QUBO problem, each variable in the QUBO formulation is mapped onto a corresponding neuron, and the pairwise interactions defined by the QUBO matrix translate directly into synaptic connections. In a manner analogous to simulated annealing, NEBM orchestrates a progressive reduction of external driving influences, guiding the network toward low-energy states. The non-equilibrium nature of the model ensures that the system remains driven by both internal neuronal processes and external inputs throughout its evolution. Ultimately, this paradigm is well-suited for deployment on neuromorphic platforms such as Loihi 2, where spiking neural networks can efficiently implement these dynamic, non-equilibrium processes to solve complex optimization problems.
2.3.1 State Variables and Parameters
Each NEBM neuron $i$ maintains at least the following state variables:
  • $v_{i}(t)$: The membrane potential at time step $t$.
  • $a_{i}(t)$: The binary $\{0,1\}$ spike output at time $t$.
  • $T(t)$: The system temperature (or inverse temperature $\beta(t)\equiv 1/T(t)$) follows an annealing schedule.
  • $r_{i}(t)$: A refractory state or counter to modulate firing behavior over time.
  • $\alpha,\rho,\kappa,\gamma,\theta$: Hyperparameters controlling decay factors, refractory penalties, noise scales, and thresholds.
2.3.2 Membrane Potential Update
The discrete-time update for the membrane potential is:
$v_{i}(t+1)\;=\;\alpha\,v_{i}(t)\;+\;\sum_{j}Q_{ij}\,a_{j}(t)\;+\;h_{i}\;+\;\gamma\,\eta_{i}\bigl{(}t;T(t)\bigr{)},$ (6)
where:
  • $\alpha\in[0,1)$ is a decay constant ensuring $v_{i}$ does not grow unbounded.
  • $Q_{ij}$ is the weight from neuron $j$ to $i$, derived from the QUBO matrix $Q$.
  • $h_{i}$ is the bias term for neuron $i$, from the vector $\mathbf{h}$.
  • $\gamma\,\eta_{i}(t;T)$ represents noise injection of a particular distribution, scaled by $\gamma$. The temperature $T(t)$ typically controls the noise variance or amplitude.
2.3.3 Spiking / Stochastic “Boltzmann-like” Activation
After updating the membrane potential, the neuron generates an output spike $s_{i}(t+1)$. We denote $v_{i}(t+1)$ as the updated membrane potential for neuron $i$ at time $t+1$. Although a deterministic threshold is available, we chose a biologically inspired and exploration-friendly approach to compute the spike output stochastically:
$\mathbb{P}\bigl{(}a_{i}(t+1)=1\bigr{)}\;=\;\sigma\!\Bigl{(}\frac{v_{i}(t+1)}{T(t)}\Bigr{)}\;=\;\frac{1}{1+e^{-\frac{v_{i}(t+1)}{T(t)}}},$ (7)
where $\sigma(\cdot)$ is the logistic function, and $T(t)$ is the temperature. The higher the membrane potential (relative to temperature), the greater the probability of spiking.
2.3.4 Refractory and Reset Dynamics
A refractory mechanism helps mimic annealing and prevents excessive firing. Our approach is:
$r_{i}(t+1)\;=\;\rho\,r_{i}(t)\;+\;a_{i}(t),$ (8)
$v_{i}(t+1)\;=\;\alpha\,v_{i}(t)\;+\;\sum_{j}Q_{ij}\,a_{j}(t)\;+\;h_{i}\;-\;\kappa\,r_{i}(t)\;+\;\gamma\,\eta_{i}\bigl{(}t;T(t)\bigr{)},$ (9)
where:
  • $\rho\in(0,1)$ is a decay factor for the refractory state $r_{i}$.
  • $\kappa$ is a refractory scaling term that reduces the membrane potential based on recent spiking.
If $r_{i}(t)$ exceeds a certain threshold, the neuron can also be prevented from firing.
2.3.5 Annealing Schedule
To help the network converge on low-energy solutions, the temperature $T$ decreases over time:
$T(t)\;=\;T_{\max}\;-\;\Delta_{T}\,\cdot t,$ (10)
where $T_{\max}$ is the initial temperature and $\Delta_{T}$ is a constant decrement (for a linear schedule). As $T$ lowers, the system becomes more likely to settle into low-energy (potentially optimal) configurations rather than continuing to hop among many states.
2.3.6 Implementation on Loihi